On January 1, 2018, Ameen Company purchased major pieces of manufacturing equipment for a total of $114 million. Ameen uses straight-line depreciation for financial statement reporting and deducted 100% of the equipment’s cost for income tax reporting in 2018. At December 31, 2020, the book value of the equipment was $108 million. At December 31, 2021, the book value of the equipment was $80 million. There were no other temporary differences and no permanent differences. Pretax accounting income for 2021 was $128 million.

Required:
1. Prepare the appropriate journal entry to record Ameen’s 2021 income taxes. Assume an income tax rate of 25%.
2. What is Ameen’s 2021 net income?
Answer #1

Ans 1:

Taxable income for 2021 = Pretax accounting income + Excess of book depreciation over tax depreciation

= $128 + [($108 - $80) - 0]

= $156 million

Income tax payable = $156*25% = $39 million

Reversal of deferred tax liability = $28*25% = $7million

*108-80=28

Journal Entries - Ameen Company (In millions)

Date

Particulars

Debit

Credit

31-Dec-21

Income tax expense Dr

$32.00

Deferred tax liability Dr

$7.00

       To Income taxes Payable ($156*25%)

$39.00

(To record income tax and deferred tax for 2021)

Ans 2:

Net income for 2021 = Pretax income - Income tax expense = $128 - $32 = $96 million